U/I Question 2 of 17 How many atoms are in 153 g of calcium? 28 23.0 x10 153 g
Answer #1

Molar mass of Ca = 40.08 g/mol

mass(Ca)= 153 g

use:

number of mol of Ca,

n = mass of Ca/molar mass of Ca

=(1.53*10^2 g)/(40.08 g/mol)

= 3.817 mol

use:

number of atoms = number of moles * Avogadro’s number

= 3.817 * 6.022*10^23 atoms

= 2.30*10^24 atoms

Answer: 2.30*10^24 atoms